OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 2. Find x when y = 4

OpenStudy (anonymous):

y=k/3x 2=k/3*6 k=2*18*36 y=36/3x=12/x 4=12/x => x=12/4 = 3

OpenStudy (anonymous):

y=3x x=3y x/3=y x/3 =4 x=12